A shot putter exerts a force of 100N acting forward and upward on a 40N shot...

  1. A shot putter exerts a force of 100N acting forward and upward on a 40N shot put at an angle of 60° from horizontal. Other than gravity, the weight of the shot put (40N) is the only external force acting on it.

a. What was the vertical component of this force?

b. What is the horizontal component of this force?

c. How big was the NET force (sum of all external forces) acting on the shot put? (remember gravity is also acting in the horizontal direction)
Answer #1

a)

Fx = 100 cos 60 = 50 N

b)

Fy = 100 sin 60 = 86.6 N

c)net vertical force

Fv = 86.6 - 40 = 46.6 N

net force

F^2 = Fv^2 + Fx^2

F = 68.3505 N

======

Comment before rate in case any doubt, will reply for sure.. goodluck